# ASGO: Adaptive Structured Gradient Optimization

### v2

@@ Abstract
Training deep neural networks is a structured optimization problem, because the parameters are naturally represented by matrices and tensors rather than by vectors. Under this structural representation, it has been widely observed that gradients are low-rank and Hessians are approximately block-wise diagonal. These structured properties are crucial for designing efficient optimization algorithms, but are not utilized by many current popular optimizers like Adam. In this paper, we present a novel optimization algorithm ASGO that capitalizes on these properties by employing a preconditioner that is adaptively updated using structured gradients. By fine-grained theoretical analysis, ASGO is proven to achieve superior convergence rates compared to existing structured gradient methods. Based on the convergence theory, we further demonstrate that ASGO can benefit from the low-rank and block-wise diagonal properties. We also discuss practical modifications of ASGO and empirically verify ASGO’s effectiveness on language model tasks.
@@ 1 Introduction
Numerical optimization algorithms, especially those can efficiently train large foundation models (Devlin et al., 2018; Brown et al., 2020; Touvron et al., 2023, ; Ouyang et al., 2022), play an important role in the modern machine learning field. Among them, adaptive gradient methods like AdaGrad (Duchi et al., 2011) and Adam (Kingma and Ba, 2014) are popular choices, gaining huge success in training state-of-the-art models in many tasks. These algorithms typically apply a diagonal matrix preconditioner to the gradient $g_{t}$ to update the deep neural network (DNN) parameters $w_{t}$;
$w_{t+1}=w_{t}-\eta_{t}\Lambda_{t}^{-1}g_{t},\text{ where $w_{t}\in\mathbb{R}^{d}$, $g_{t}\in\mathbb{R}^{d}$, and $\Lambda_{t}\in\mathbb{R}^{d\times d}$ is a diagonal matrix}.$
This coordinate-wise step size design has been theoretically verified to be effective as it can exploit the sparsity of the gradient vectors $g_{t}$ (Duchi et al., 2011). Also, when the Hessian is well-approximated by a diagonal matrix whose diagonal entries have very different scales, adaptive gradient methods have been proven to be beneficial (Liu et al., 2024; Jiang et al., 2024a; Xie et al., 2024). While these results seem to be convincing, common DNNs do not necessarily have sparse gradients or Hessians that are well-approximated by ill-conditioned diagonal matrices. Instead, if we take the matrix structure of gradients in neural networks into account, it has been widely observed that these structured gradients are usually low-rank (Zhao et al., 2021; Yang et al., 2023; Cosson et al., 2023), and the Hessians are well-approximated by block-wise diagonal matrices (Collobert, 2004; Zhang et al., 2024a, b). Since adaptive gradient methods like Adam, treat the parameters as vectors and ignore the matrix structure of gradients, they are generally unable to exploit these structured properties. This gap makes us ask the following question: How can we properly consider the matrix structures of gradients and exploit their low-rank and block-wise diagonal properties?
@@
$W_{t+1}=W_{t}-\eta_{t}L_{t}^{-\frac{1}{4}}G_{t}R_{t}^{-\frac{1}{4}},\text{ where $W_{t},G_{t}\in\mathbb{R}^{m\times n}$ and $L_{t}\in\mathbb{R}^{m\times m}$, $R_{t}\in\mathbb{R}^{n\times n}$ are full matrices.}$.
@@
In this paper, we provide an answer to the aforementioned question by proposing ASGO (Adaptive Structured Gradient Optimization), which significantly improves the convergence guarantees of Shampoo while requiring less memory and computation. In light of the analysis that demonstrates the benefits of adaptive gradient methods (Duchi et al., 2011; Liu et al., 2024; Jiang et al., 2024a; Xie et al., 2020), we use appropriate assumptions to enable a more fine-grained convergence analysis, showing superior convergence results and how ASGO can benefit from the low-rank and block-wise diagonal properties of the problem. We also discuss the connection between ASGO and Muon (Jordan et al., 2024), a structured gradient method based upon the steepest descent algorithm with spectral norm, to conjecture a relation between ASGO and Muon analogous to the relation between AdaGrad and SignSGD (Bernstein et al., 2018; Kunstner et al., 2023). Furthermore, we develop an efficient design for query-key attention parameters in transformer models and examine the empirical performance on pretraining transformer model tasks. Our main contributions are summarized as follows.
@@
  • We further demonstrate that ASGO can effectively exploit the low-rankness of gradients as well as the approximate block-wise diagonal property of Hessians that is typically observed in training DNNS, hence indicating great potential of ASGO for real-world applications.
  • We develop a practical implementation of ASGO with targeted modifications for transformer architectures that offers significant advantages: it eliminates the need for a separate optimizer for 1D parameters in contrast to Muon, while requiring less memory and computational complexity than Shampoo.
  • We empirically validate the effectiveness of ASGO with modifications on language model tasks, demonstrating the algorithm’s great potential in real applications.
@@ 2 Related Work
Adaptive gradient methods that use diagonal preconditioners to speedup the convergence are extremely popular for solving many real-world optimization problems. To the best of our knowledge, the first method of this kind for machine learning, AdaGrad (Duchi et al., 2011; Streeter and McMahan, 2010), was developed based on rigorous theory that showed the benefits of using this kind of preconditioner. Adam (Kingma and Ba, 2014; Loshchilov and Hutter, 2017) modified AdaGrad and has become the default choice for training large foundation models. In theory, it has been proven that adaptive gradient methods can benefit from sparse gradients and approximately ill-conditioned Hessians (Duchi et al., 2011; Liu et al., 2024; Jiang et al., 2024a; Xie et al., 2024). It is worth noting that the original AdaGrad paper (Duchi et al., 2011) also proposed a version of AdaGrad that uses a full-matrix preconditioner instead of the diagonal one, which is believed to perform even better. However, this full-matrix AdaGrad method suffers from large memory costs for storing the preconditioner, and there is no better convergence guarantee compared to diagonal AdaGrad and SGD under the same settings as listed in Section 4.
@@
In the standard optimization literature (i.e., excluding areas such as conic optimization), it is common to consider variables as vectors. However, recently, optimization methods for machine learning that consider variables as matrices have been rapidly gaining attention. Adafactor (Shazeer and Stern, 2018), LAMB (You et al., 2019), and Adam-mini (Zhang et al., 2024b) consider the matrix or layer structure to help reduce the memory cost of Adam and enable more efficient training. Shampoo (Gupta et al., 2018) and KFAC (Martens and Grosse, 2015) are two pioneering works that approximate, respectively, the full-matrix preconditioner of AdaGrad and the true Fisher matrix (FM), using Kronecker products of smaller matrices, making the memory cost more affordable. Unfortunately, the analysis in (Gupta et al., 2018) shows that the rate of convergence for Shampoo is no better than that for the full-matrix AdaGrad. We note that there is another method, TNT (Ren and Goldfarb, 2021) that is very closely related to Shampoo. TNT was developed as a natural gradient method, approximating the true FM by the covariance of block-wise sampling-based gradients assuming that they are Tensor-Normally distributed. The main difference between TNT and Shampoo, is that TNT uses true FMs and inverses of their Kronecker factors, whereas Shampoo uses empirical FMs and the -1/4 power of their factors. More discussion on recent and concurrent work is presented in Appendix A.
@@ 3.1 Notation and problem setting
Throughout this paper, we use capital letters such as $W$ to represent matrices and $[W]_{i,j}$ to denote the $(i,j)$-th entry of $W$. For an arbitrary matrix $W\in\mathbb{R}^{m\times n}$, we denote
  • $\left\|W\right\|_{\mathrm{op}}$ as the spectral norm of a matrix $W$, i.e., the largest singular value of it;
@@
For symmetric square matrices, $A,B\in\mathbb{R}^{m\times m}$, $A\preceq B$ denotes that $B-A$ is positive semidefinite and $A\prec B$ denotes that $B-A$ is positive definite. $\succ$ and $\succeq$ are defined accordingly. We study the following stochastic optimization problem:
@@ 3.2 ASGO (Algorithm 1)
$W_{t+1}=W_{t}-\eta_{t}V_{t}^{-\frac{1}{2}}G_{t},\text{ where $W_{t}\in\mathbb{R}^{m\times n}$ and $V_{t}\in\mathbb{R}^{m\times m}$ is a full matrix.}$.
ASGO only needs to store one preconditioner matrix and compute its matrix square root and inverse on each iteration, versus two matrices for Shampoo. On the other hand, ASGO’s preconditioner may not be as good an approximation to the full-matrix AdaGrad preconditioner or the empirical Fisher matrix (Gupta et al., 2018; Morwani et al., 2024) as Shampoo’s. However, as we shall see in the following sections, this design can actually better exploit the low-rankness of gradients and the block-wise diagonal nature of Hessians, in terms of achieving better convergence rates.
@@ 4 Nonsmooth Theory
Note that we treat $D_{\mathrm{op}}$ as a constant here, but it may increase as the iteration number $T$ increases. This can be addressed, for example, by invoking a projection onto a bounded convex set $\mathcal{W}$ with respect to the norm $\left\|\cdot\right\|_{\Lambda_{t}}$ in each iteration, as in AdaGrad (Duchi et al., 2011). $D_{\mathrm{op}}$ would then be bounded by the spectral norm of $\mathcal{W}$. However, since this projection is rarely used in training DNNs, we follow Gupta et al. (2018) and omit it in Algorithm 1. More importantly, this theoretical bound depends on the trace norm of gradients and the spectral norm of weights, showing that the algorithm can make use of the low-rank property of gradients.
One can easily check that this convergence rate for convex nonsmooth problems is $\mathcal{O}(1/\sqrt{T})$, the same as SGD (Zinkevich, 2003) (see below) and AdaGrad (Duchi et al., 2011). Comparison with SGD. The convergence rate for SGD under the assumptions of Corollary 2 is:
[…]
@@
  • $\left\|Q\right\|_{\mathrm{F}}\leq\left\|Q\right\|_{*}\leq\sqrt{r_{G}}\left\|Q\right\|_{\mathrm{F}}$, where $r_{G}$ is the rank of $Q$. Thus when $G_{t}$ are low-rank, or have very imbalanced singular values, $\left\|Q\right\|_{*}$ can be close to $\left\|Q\right\|_{\mathrm{F}}$;
  • $D_{\mathrm{F}}/\sqrt{r_{D}}\leq D_{\mathrm{op}}\leq D_{\mathrm{F}}$, where $r_{D}=\text{max rank of}(W_{t}-W_{*})$. Thus, when $W_{t}-W_{*}$ are relatively high-rank or have lots of singular values of a similar scale, $D_{\mathrm{op}}$ can be much smaller than $D_{\mathrm{F}}$.
Therefore, ASGO should work well when $G_{t}$ are low-rank and $W_{t}-W_{*}$ are relatively high-rank. Intuitions on ASGO in practical tasks. We argue that low-rank $G_{t}$ and relatively high-rank $W_{t}-W_{*}$ should be common in many practical tasks, revealing the potential of ASGO in applications. As we have discussed in Section 2, gradients are commonly low-rank in DNNs, as verified by Zhao et al. (2021); Yang et al. (2023); Cosson et al. (2023). Meanwhile, given the huge success of LoRA (Hu et al., 2022) in fine-tuning foundation models, which naturally asserts a low-rank total update in $W$, there may seem to be a conflict to assume that $W_{t}-W_{*}$ has a high rank. However, it has been observed that $W_{0}-W_{*}$ should be relatively high-rank to obtain a better result, at least in pretraining and some complex fine-tuning tasks for large foundation models (Lialin et al., 2023; Jiang et al., 2024b; Huang et al., 2025). Further exploration of the connection between the rank of weight updates and the performance of algorithms is an interesting topic for future research. Comparison with Full-Matrix AdaGrad and Shampoo. Shampoo (Gupta et al., 2018) and the full-matrix AdaGrad (Duchi et al., 2011) achieve the following convergence rates under the same settings as Theorem 1.
[…]
@@
Proofs of all of the results in this section are given in Appendix E. We keep the matrix structure of $W_{t}$ and $G_{t}$ throughout our analysis, in contrast to standard analyses of the convergence of AdaGrad and Shampoo, which are based on vectorizations of $W_{t}$ and $G_{t}$. This is important for proving that ASGO can exploit the structured properties.
@@ 5 Smooth Theory
If $X=[x_{1},\dots x_{n}]$, where each $x_{i}\in\mathbb{R}^{m}$, and we vectorize $X$, i.e., $\mathbf{x}\equiv\mathrm{vec}(X)=(x_{1}^{T},\ldots,x_{n}^{T})^{T}$, and form the block diagonal matrix $\mathbf{L\equiv}\mathrm{diag}[L,\dots,L]$, we obtain that ${\rm tr}\left(X^{\top}LX\right)=\sum_{i=1}^{n}x_{i}^{\top}Lx_{i}=\mathbf{x}^{T}\mathbf{L}\mathbf{x}.$ This means that Assumption 2 is equivalent to the existence of a symmetric matrix $L\in\mathbb{R}^{m\times m},L\succ 0$ such that for any $w\in\mathbb{R}^{mn}$, $-\mathbf{L}\preceq\nabla^{2}f_{v}(\mathbf{w})\preceq\mathbf{L},$ where $f_{v}(\mathbf{w})=f(W)$, $W=[w_{1},\dots,w_{n}]\in\mathbb{R}^{m\times n}$ and $\mathbf{w}\equiv\mathrm{vec}(W)=[w_{1}^{T},\dots,w_{n}^{T}]^{T}\in\mathbb{R}^{mn}$. Hence, it is closely related to the block-wise diagonal structure of the Hessian as observed by many researchers; (see Figure 3 in Appendix A). Also note that Assumption 2 implies the standard smoothness assumption with respect to the Frobenius norm by $\left\|\nabla f_{v}(\mathbf{w})\right\|_{\mathrm{op}}\leq\left\|L\right\|_{\mathrm{op}}$. This block-wise diagonal smoothness is an extension of the standard smoothness, which is related to but different from the diagonal anisotropic smoothness employed for analyzing sign-based and adaptive gradient methods (Bernstein et al., 2018; Liu et al., 2024).
@@
One can check that Assumption 3 implies the standard variance bound $\mathbb{E}[\left\|N_{t}\right\|_{\mathrm{F}}^{2}]\leq\left\|V\right\|_{\mathrm{F}}^{2}$. The assumption shares some similarity with the coordinate-wise variance bounds in Bernstein et al. (2018); Crawshaw et al. (2022); Liu et al. (2024), in the sense that it allows a more fine-grained analysis. This matrix-form variance upper bound may better describe the real case since it takes the structure of the noise into account, which is relevant to matrix rank and other structured properties.
@@
[…]
where $D_{\mathrm{F}}$ is defined in Theorem 3. Since the comparison between the $\mathcal{O}(1/\sqrt{MT})$ term is generally consistent with the discussion in Section 4, we further compare the $\mathcal{O}(1/T)$ term here to provide some more intuition: we have
@@
Therefore, we can see that in general, ASGO should work well when the Hessian can be well approximated by a block-wise diagonal matrix with a low-rank $L$ for each block and $W_{t}-W_{*}$ are of relatively high-rank. Note that Hessians have been found to be low-rank in DNNs, especially after some steps of training (Sagun et al., 2016, 2017; Wu et al., 2020). Block-wise diagonal structure of DNNs. It has been widely observed that the Hessian of MLPs and other DNNs, are approximately block-wise diagonal (Collobert, 2004; Zhang et al., 2024a, b; Bahamou et al., 2022), as shown in Figure 3 in Appendix A. This block-wise diagonal structure naturally arises from the structure of MLPs, where all the parameters associated with a particular neuron (i.e., all elements of either a row or column of $W$) are more closely related and have a denser block Hessian matrix, than a set of parameters that do not share such an association. Intuitively, the elements of the rows of $W$ seem to be more likely to be closely related to the columns of $W$. Based on Theorem 3, our algorithm should perform well in this block Hessian setting, showing great potential in real applications. Intuition on single-side preconditioners. Since ASGO only uses single-side preconditioners, it is important to determine on which side they should be applied. As demonstrated in Figure 3, a block in the Hessian commonly corresponds to all input weights into a neuron in the next layer, i.e., as a row of $W$. Therefore, we may want the single-side preconditioner to be on the right side of $G_{t}$. Based on this block-wise diagonal structure, Shampoo-like double-sided preconditioners may not only suffer from higher memory and computation costs, but also may precondition less effectively, since they are trying to approximate the structured curvature information matrix using a dense matrix produced by a Kronecker product (Gupta et al., 2018; Morwani et al., 2024).

### v3

@@ Abstract
Training deep neural networks is a structured optimization problem, because the parameters are naturally represented by matrices and tensors rather than by vectors. Under this structural representation, it has been widely observed that gradients are low-rank and Hessians are approximately block diagonal. These structured properties are crucial for designing efficient optimization algorithms, but are not utilized by many current popular optimizers like Adam. In this paper, we present a novel optimization algorithm ASGO that capitalizes on these properties by employing a preconditioner that is adaptively updated using structured gradients. By a fine-grained theoretical analysis, ASGO is proven to achieve superior convergence rates compared to existing structured gradient methods. Based on this convergence theory, we further demonstrate that ASGO can benefit from low-rank gradients and block diagonal Hessians. We also discuss practical modifications of ASGO and empirically verify ASGO’s effectiveness on language model tasks. Code is available at https://github.com/infinity-stars/ASGO.
@@ 1 Introduction
Numerical optimization algorithms, especially those that can efficiently train large foundation models (Devlin et al., 2018; Brown et al., 2020; Touvron et al., 2023a, b; Ouyang et al., 2022), play an important role in the modern machine learning field. Among them, adaptive gradient methods like AdaGrad (Duchi et al., 2011) and Adam (Kingma and Ba, 2014) are popular choices, gaining huge success in training state-of-the-art models in many tasks. These algorithms typically apply a diagonal matrix preconditioner to the gradient $g_{t}$ to update the deep neural network (DNN) parameters $w_{t}$;
[…]
This coordinate-wise step size design has been theoretically verified to be effective as it can exploit the sparsity of the gradient vectors $g_{t}$ (Duchi et al., 2011). Also, when the Hessian is well-approximated by a diagonal matrix whose diagonal entries have very different scales, adaptive gradient methods have been proven to be beneficial (Liu et al., 2024; Jiang et al., 2024a; Xie et al., 2024). While these results seem to be convincing, common DNNs do not necessarily have sparse gradients or Hessians that are well-approximated by ill-conditioned diagonal matrices. Instead, if we take the matrix structure of gradients in DNNs into account, it has been widely observed that these structured gradients are usually low-rank (Zhao et al., 2021; Yang et al., 2023; Cosson et al., 2023), and the Hessians are well-approximated by block diagonal matrices (Collobert, 2004; Zhang et al., 2024a, b). Since adaptive gradient methods like Adam, treat the parameters as vectors and ignore the matrix structure of gradients, they are generally unable to exploit these structured properties. This leads us to ask the following question: How can we effectively use the matrix structure of gradients to exploit their low-rank and block diagonal properties?
@@
$W_{t+1}=W_{t}-\eta_{t}L_{t}^{-\frac{1}{4}}G_{t}R_{t}^{-\frac{1}{4}},\text{ where $W_{t},G_{t}\in\mathbb{R}^{m\times n}$ and $L_{t}\in\mathbb{R}^{m\times m}$, $R_{t}\in\mathbb{R}^{n\times n}$ are full matrices.}$
@@
In this paper, we provide an answer to the aforementioned question by proposing ASGO (Adaptive Structured Gradient Optimization), which significantly improves the convergence guarantees of Shampoo while requiring less memory and computation. In light of the analysis that demonstrates the benefits of adaptive gradient methods (Duchi et al., 2011; Liu et al., 2024; Jiang et al., 2024a; Xie et al., 2020), we use appropriate assumptions to enable a more fine-grained convergence analysis, that yields superior convergence results and an explanation of how ASGO can benefit from the low-rank and block diagonal properties of a given problem. We also discuss the connection between ASGO and Muon (Jordan et al., 2024), a structured gradient method based upon the steepest descent algorithm with respect to the spectral norm, to conjecture a relation between ASGO and Muon analogous to the relation between AdaGrad and SignSGD (Bernstein et al., 2018; Kunstner et al., 2023). Furthermore, we develop an efficient design for query-key attention parameters in transformer models and examine its empirical performance on pretraining transformer model tasks. Our main contributions are summarized as follows.
@@
  • We further demonstrate that ASGO can effectively exploit the low-rankness of gradients as well as the approximate block diagonal property of Hessians that is typically observed in training DNNS, indicating good theoretical properties under realistic settings.
  • We develop a practical implementation of ASGO that is efficient while enhancing ASGO’s effectiveness. We further empirically validate the effectiveness and efficiency of this implementation on language model tasks, demonstrating the algorithm’s great potential in real applications.
@@ 2 Related Work
Adaptive gradient methods that use diagonal preconditioners to speedup convergence are extremely popular for solving many real-world optimization problems. To the best of our knowledge, the first method of this kind for machine learning, AdaGrad (Duchi et al., 2011; Streeter and McMahan, 2010), was developed based on rigorous theory that showed the benefits of using this kind of preconditioner. Adam (Kingma and Ba, 2014; Loshchilov and Hutter, 2017) modified AdaGrad and has become the default choice for training large foundation models. In theory, it has been proven that adaptive gradient methods can benefit from sparse gradients and approximately ill-conditioned Hessians (Duchi et al., 2011; Liu et al., 2024; Jiang et al., 2024a; Xie et al., 2024). It is worth noting that the original AdaGrad paper (Duchi et al., 2011) also proposed a version of AdaGrad that uses a full-matrix preconditioner instead of the diagonal one, which is believed to perform even better. However, this full-matrix AdaGrad method suffers from large memory costs for storing the preconditioner, while having no better convergence guarantee compared to diagonal AdaGrad and SGD under the assumption of objective function convexity as described in Section 4 below.
@@
In the standard optimization literature (i.e., excluding areas such as conic optimization), it is common to consider variables as vectors. However, recently, optimization methods for machine learning that consider variables as matrices have been rapidly gaining attention. Adafactor (Shazeer and Stern, 2018), LAMB (You et al., 2019), and Adam-mini (Zhang et al., 2024b) consider the matrix or layer structure to help reduce the memory cost of Adam and enable more efficient training. Shampoo (Gupta et al., 2018) and KFAC (Martens and Grosse, 2015) are two pioneering works that approximate, respectively, the full-matrix preconditioner of AdaGrad and the true Fisher matrix (FM), using Kronecker products of smaller matrices, making the memory cost more affordable. Unfortunately, the analysis in (Gupta et al., 2018) shows that the rate of convergence for Shampoo is no better than that for the full-matrix AdaGrad. We note that there is another method, TNT (Ren and Goldfarb, 2021) that is very closely related to Shampoo. TNT was developed as a natural gradient method, approximating the true FM by the covariance of block-wise sampling-based gradients assuming that they are Tensor-Normally distributed. The main difference between TNT and Shampoo, is that TNT uses true FMs and inverses of their Kronecker factors, whereas Shampoo uses empirical FMs and the -1/4 power of their factors. More discussion on related recent and concurrent work is presented in Appendix A.
@@ 3.1 Notation and problem setting
Throughout this paper, we use capital letters such as $W$ to represent matrices and $W]_{i,j}$ to denote the $(i,j)$-th entry of $W$. For an arbitrary matrix $W\in\mathbb{R}^{m\times n}$, we denote
  • $\left\|W\right\|_{\mathrm{op}}$ as the spectral norm of a matrix $W$, i.e., its largest singular value;
@@
For symmetric square matrices, $\in\mathbb{R}^{m\times m}, ABA\preceq B$ denotes that BAB-A is positive semidefinite and $A\prec B$ denotes that BAB-A is positive definite. $\succ$ and $\succeq$ are defined accordingly. We study the following stochastic optimization problem:
@@ 3.2 ASGO (Algorithm 1)
[…]
ASGO only needs to store one preconditioner matrix and compute its matrix square root and inverse on each iteration, versus two matrices for Shampoo. On the other hand, ASGO’s preconditioner may not be as good an approximation to the full-matrix AdaGrad preconditioner or the empirical Fisher matrix (Gupta et al., 2018; Morwani et al., 2024) as Shampoo’s. However, as we shall see in the following sections, this design can exploit the low-rankness of gradients and the block diagonal nature of Hessians to achieve better convergence rates.
@@ 4 Nonsmooth Theory
Note that we treat $D_{\mathrm{op}}$ as a constant here, but it may increase as the iteration number $T$ increases. This can be addressed, for example, by invoking a projection onto a bounded convex set $\mathcal{W}$ with respect to the norm $\left\|\cdot\right\|_{\Lambda_{t}}$ in each iteration, as in AdaGrad (Duchi et al., 2011). $D_{\mathrm{op}}$ would then be bounded by the spectral norm of $\mathcal{W}$. However, since this projection is rarely used in training DNNs, we follow Gupta et al. (2018) and omit it in Algorithm 1. We refer interested readers to Appendix H for a discussion of how one can incorporate the projection operation and obtain $D_{\mathrm{op}}$ and $D_{\mathrm{F}}$ independent of $T$. More importantly, this theoretical bound depends on the trace norm of gradients and the spectral norm of weights, showing that the algorithm can make use of the low-rank property of gradients.
One can easily check that this convergence rate for convex nonsmooth problems is $\mathcal{O}(1/\sqrt{T})$, the same as SGD (Zinkevich, 2003) (see below) and AdaGrad (Duchi et al., 2011).
Comparison with SGD. The convergence rate for SGD under the assumptions of Corollary 2 is:
$\text{SGD:}\quad\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[f(W_{t})]-f(W_{*})\leq\mathcal{O}\left(\frac{D_{\mathrm{F}}\left\|Q\right\|_{\mathrm{F}}}{\sqrt{T}}\right),$
@@
  • $\left\|Q\right\|_{\mathrm{F}}\leq\left\|Q\right\|_{*}\leq\sqrt{r_{G}}\left\|Q\right\|_{\mathrm{F}}$, where $r_{G}$ is the rank of $Q$. Thus when the $G_{t}$ are low-rank, or have very imbalanced singular values, $\left\|Q\right\|_{*}$ can be close to $\left\|Q\right\|_{\mathrm{F}}$;
  • $D_{\mathrm{F}}/\sqrt{r_{D}}\leq D_{\mathrm{op}}\leq D_{\mathrm{F}}$, where $r_{D}=\text{max rank of}(W_{t}-W_{*})$. Thus, when the $W_{t}-W_{*}$ are relatively high-rank or have lots of singular values of a similar scale, $D_{\mathrm{op}}$ can be much smaller than $D_{\mathrm{F}}$.
Therefore, ASGO should work well when $G_{t}$ are low-rank and $W_{t}-W_{*}$ are relatively high-rank.
Intuition about ASGO in practice. We argue that $G_{t}$ is low-rank and $W_{t}-W_{*}$ is often relatively high-rank in many practical tasks, revealing the potential of ASGO in applications. As we have discussed in Section 2, gradients are commonly low-rank in DNNs, as verified by Zhao et al. (2021); Yang et al. (2023); Cosson et al. (2023). Meanwhile, given the huge success of LoRA (Hu et al., 2022) in fine-tuning foundation models, which employs low-rank total updates of $W$, there may seem to be a conflict to assume that $W_{t}-W_{*}$ has a high rank. However, it has been observed that $W_{0}-W_{*}$ should be relatively high-rank to obtain a better result, at least in pretraining and some complex fine-tuning tasks for large foundation models (Lialin et al., 2023; Jiang et al., 2024b; Huang et al., 2025). Further exploration of the connection between the rank of weight updates and the performance of algorithms is an interesting topic for future research.
Comparison with Full-Matrix AdaGrad and Shampoo. Shampoo and the full-matrix AdaGrad achieve the following convergence rates under the same settings as Theorem 1.
$\text{Full-Matrix AdaGrad:}\quad\mathcal{O}\left(D_{\mathrm{F}}\sum_{j=1}^{m}\sum_{i=1}^{n}\sqrt{\sum_{t=0}^{T-1}[G_{t}]_{i,j}^{2}}\right)$
$\text{Shampoo:}\quad\mathcal{O}\left(\sqrt{r}D_{\mathrm{F}}\cdot{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}G_{t}^{\top}\right)^{\frac{1}{4}}\right)\cdot{\rm tr}\left(\left(\sum_{t=0}^{T-1}G_{t}^{\top}G_{t}\right)^{\frac{1}{4}}\right)\right).$
@@
Proofs of all of the results in this section are given in Appendix E. We keep the matrix structure of $W_{t}$ and $G_{t}$ throughout our analysis, in contrast to standard analyses of the convergence of AdaGrad and Shampoo, which are based on vectorizations of $W_{t}$ and $G_{t}$. This is important for proving that ASGO can exploit the structured properties. We note that for simplicity in Algorithm 1 above, the preconditioner operates on left-side of $G_{t}$. However, in a practical implementation of ASGO, using the right-sided preconditioner is preferable unless the number of columns of $G_{t}$ is much greater than the number of rows. See the discussion toward the end of Section 5 and in Section 7.1.2 and Algorithm 2 in Appendix B.
@@ 5 Smooth Theory
If $X=[x_{1},\dots x_{n}]$, where each $x_{i}\in\mathbb{R}^{m}$, and we vectorize $X$, i.e., $\mathbf{x}\equiv\mathrm{vec}(X)=(x_{1}^{T},\ldots,x_{n}^{T})^{T}$, and form the block diagonal matrix $\mathbf{L\equiv}\mathrm{diag}[L,\dots,L]$, we obtain that ${\rm tr}\left(X^{\top}LX\right)=\sum_{i=1}^{n}x_{i}^{\top}Lx_{i}=\mathbf{x}^{T}\mathbf{L}\mathbf{x}.$ This means that Assumption 2 is equivalent to the existence of a symmetric matrix $L\in\mathbb{R}^{m\times m},L\succ 0$ such that for any $w\in\mathbb{R}^{mn}$, $\mathbf{L}\preceq\nabla^{2}f_{v}(\mathbf{w})\preceq\mathbf{L},$ where $f_{v}(\mathbf{w})=f(W)$, $W=[w_{1},\dots,w_{n}]\in\mathbb{R}^{m\times n}$ and $\mathbf{w}\equiv\mathrm{vec}(W)=[w_{1}^{T},\dots,w_{n}^{T}]^{T}\in\mathbb{R}^{mn}$. Hence, it is closely related to the block-wise diagonal structure of the Hessian as observed by many researchers with each block corresponding to a column of $W$; (see Figure 3 in Appendix A). Also note that Assumption 2 implies the standard smoothness assumption with respect to the Frobenius norm by $\left\|\nabla f_{v}(\mathbf{w})\right\|_{\mathrm{op}}\leq\left\|L\right\|_{\mathrm{op}}$. This block-wise diagonal smoothness is an extension of the standard smoothness, which is related to but different from the diagonal anisotropic smoothness employed for analyzing sign-based and adaptive gradient methods (Bernstein et al., 2018; Liu et al., 2024).
@@
One can check that Assumption 3 implies the standard variance bound $\mathbb{E}[\left\|N_{t}\right\|_{\mathrm{F}}^{2}]\leq\left\|V\right\|_{\mathrm{F}}^{2}$. The assumption shares some similarity with the coordinate-wise variance bounds in Bernstein et al. (2018); Crawshaw et al. (2022); Liu et al. (2024), in the sense that it allows a more fine-grained analysis. This matrix-form variance upper bound may better describe the real case since it takes the structure of the noise into account, which is relevant to matrix rank and other structural properties.
@@
$\text{SGD:}\quad\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[f(W_{t})]-f(W_{*})\leq\mathcal{O}\left(\frac{D_{\mathrm{F}}^{2}\left\|L\right\|_{\mathrm{op}}}{T}+\frac{D_{\mathrm{F}}\left\|V\right\|_{\mathrm{F}}}{\sqrt{MT}}\right),$
where $D_{\mathrm{F}}$ is defined in Theorem 3. Since the comparison between the $\mathcal{O}(1/\sqrt{MT})$ term is generally consistent with the discussion in Section 4, we further compare the $\mathcal{O}(1/T)$ term here to provide some more intuition. Specifically, we have
@@
Therefore, we can see that in general, ASGO should work well when the Hessian can be well approximated by a block diagonal matrix with a low-rank $L$ for each block and $W_{t}-W_{*}$ are of relatively high-rank. Note that Hessians have been found to be low-rank in DNNs, especially after some steps of training (Sagun et al., 2016, 2017; Wu et al., 2020).
Block diagonal structure of DNNs. It has been widely observed that the Hessian of MLPs and other DNNs, are approximately block-wise diagonal (Collobert, 2004; Zhang et al., 2024a, b; Bahamou et al., 2022), as shown in Figure 2 in Appendix A. This block-wise diagonal structure naturally arises from the structure of MLPs, where all the parameters associated with a particular neuron (i.e., all elements of a row of $W$) are more closely related and whose corresponding sub-block of the Hessian matrix is denser and has elements that are larger in absolute value than a set of parameters that do not share such an association. Intuitively, the elements of the rows of $W$ seem to be more likely to be closely related than those of the columns of $W$. Based on Theorem 3, our algorithm should perform well in this block Hessian setting, showing great potential in real applications.
Intuition on single-side preconditioners. Since ASGO only uses a single-side preconditioner, it is important to determine on which side they should be applied. As demonstrated in Figure 2, a block in the Hessian commonly corresponds to all input weights into a neuron in the next layer, i.e., as a row of the weight matrix $W$. Therefore, based on the block-wise smoothness convergence results, we should put the preconditioner $\Lambda_{t}$ on the right-hand side of $G_{t}$, i.e., computing $\Lambda_{t}=V_{t}^{1/2}$, where $V_{t}=V_{t-1}+G_{t}^{\top}G_{t}\in\mathbb{R}^{n\times n}$. In this way, we can better exploit the superior adaptability of block diagonal Hessians with correct block partitions (each row of $W$ as a block). One may also refer to Algorithm 2 for more details of our implementation. We also provide empirical results that show that right-sided outperforms left-sided preconditioning in Section 7.1.2.